OpenStudy (anonymous):

What is the function rule for the graph shown? A graph of the ordered pairs: 9 comma zero, eleven comma 2, twelve comma 3, and fourteen comma 5.

OpenStudy (anonymous):

can you write the number out

OpenStudy (anonymous):

I will just attatch the graph

OpenStudy (anonymous):

alright, let's pick out two points and find a slope betweent them

OpenStudy (anonymous):

is that picture you? cuz if it is---- HOLY CRAB APPLES YOUR A BIG MAN

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

11,2 and 12,3

OpenStudy (anonymous):

(11,2) and (12,3) slope y2 - y1 ------------------- x2- x1

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

so the second pair is x2,y2 3 - 2 ----------------- =1 12-11 that's your slope y=x+b

OpenStudy (anonymous):

y=mx+b slope=m=1 so we plug in 1 y=x+b

OpenStudy (anonymous):

so y = x + 9?

OpenStudy (anonymous):

no, 9 is x intercept, you want the y interecept

OpenStudy (anonymous):

to find b , simply plug x and y value from one ordered pair

OpenStudy (anonymous):

let's choose (11,2) y=x+b 2=11+b b=-9

OpenStudy (anonymous):

well now I am confused.. there are 4 answers and only on one of them is something other than 9.... and it is y = 1/9 * x

OpenStudy (anonymous):

wait so y = x – 9?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oooohhhhhhhhh